Find the equation of the normal of the paraboloid 3x²+4z²+4y=0 at the point (2,-4,1). Also find the point where this line again intersects the paraboloid.
1
Expert's answer
2019-07-17T11:38:54-0400

F(x,y,z) = 3x²+4y+4z2 = 0 is a paraboloid;

Fx = 6x, Fy = 4, Fz = 8z are partial derivatives.

M=(2,-4,1) is the point.


The equation of normal is

(x-x0)/Fx(M) = (y-y0)/Fy(M) = (z-z0)/Fz(M).


In our case,

(x-2)/(6*2) = (y+4)/4 = (z-1)/8


(x-2)/12 = (y+4)/4 = (z-1)/8.


In order to find the other point of intersection we need to create a parameter t:

(x-2)/12 = (y+4)/4 = (z-1)/8 = t


From that equation we get

x = 12t+2, y = 4t-4, z = 8t+1.


We put these values into the paraboloid's equation and get


3(12t+2)2+4(8t+1)2+4(4t-4) =

3(144t2+48t+4)+256t2+64t+4+16t-16=432t2+144t+12+256t2+64t+4+16t-16 =

688*t2+208t = 0.


t(688t+208) = 0.


So either t=0 (and in this case the point would be the one we already have), either t = -208/688 = -13/43.


For t = -13/43 we have:


x = 12*(-13/43)+2 = -156/43+2 = -70/43;

y = 4*(-13/43)-4 = 4*(-13/43-1) = 4*(-56/43) = -224/43;

z = 8*(-13/43)+1 = -104/43+1 = -61/43.


Answer: (-70/43, -224/43, -61/43)